OpenStudy (anonymous):

The equation of line AB is y = 5x + 1. Write an equation of a line parallel to line AB in slope-intercept form that contains point (4, 5).

OpenStudy (jmark):

Given the equation of line AB is y = 5x+1 the slope m = 1 line of equation (y-y1) = m (x-x1) substitute the value of x and y coordinate in the equation. (y-5) = 1 (x-4) y-5 = x-4 y-1 = x y = x+1 The equation of a line parallel to line AB is y = x+1. ICSE question papers for class 9 http://icse.edurite.com/icse-question-papers/icse-question-papers-for-class-9.html